OpenStudy (anonymous):

Verify the identity. cosine of x divided by quantity one plus sine of x plus quantity one plus sine of x divided by cosine of x equals two times secant of x.

OpenStudy (anonymous):

(cos x/ 1 + sin x) + (1 + sin x/ cos x) = 2 sec x That may help make a little more sense of the function.

OpenStudy (aravindg):

Try taking the LCM.

OpenStudy (anonymous):

I tried but I don't know if i am exactly doing it right.

OpenStudy (anonymous):

I got: (Cos^2 x/ Cos x + Cos x * Sin x ) + (Sin^2 + 2 sin x + 1/Cos x + Cos x * Sin x) But i don't believe i did it right.

OpenStudy (aravindg):

You are right now sin^2x+cos^2 x=?

OpenStudy (anonymous):

1?

OpenStudy (aravindg):

yeah replace sin^2x+cos^2x by 1 on numerator.

OpenStudy (anonymous):

So: 1 + 2 Sin x + 1/ Cos x + Cos x * Sin x?

OpenStudy (aravindg):

yes Now take 2 common on mumerator and cos x common on denominator.

OpenStudy (aravindg):

*numerator

OpenStudy (anonymous):

So 2 + 2 Sin x/ 2 Cos x * Sin x?

OpenStudy (aravindg):

Cos x + Cos x * Sin x=cos x(1+sin x) 2+2sin x=2(1+sin x) When you divide them 1+sin x gets cancelled. You are left with 2/cos x ie 2 sec x.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (aravindg):

yw :)